OpenStudy (anonymous):

a nut wholesaler sells a mix of peanuts and cashews. the wholesaler charges $2.65 per pound for peanuts and $4.90 per pound for cashews. The mix is to sell for $3.28 per pound. How many pounds of peanuts and how many pounds of cashews should be used to make 125 pounds of the mix?

OpenStudy (ddcamp):

Let the variable c be the total number of cahsews, and p be the number of peanuts: c + p = 125 (the total weight) 2.65×p + 4.90×c = the total cost Can you find the total cost and solve the system?